OpenStudy (anonymous):

Solve using the square root method. x2 = 64

terenzreignz (terenzreignz):

You just extract the square root of both sides?

OpenStudy (anonymous):

\[\sqrt{x ^{2}}=\sqrt{64}\]

terenzreignz (terenzreignz):

From there should be simple enough, right?

OpenStudy (anonymous):

\[x^2=64\implies x=8\text{ or } x=-8\]